OpenStudy (anonymous):

Compute the limit as X goes to zero for the function: (cosX/cos 2X)^ 1/X^2

OpenStudy (anonymous):

\[\lim_{x \to 0}\left(\frac{\cos(x)}{\cos(2x)}\right)^{\frac{1}{x^2}}\] like that?

OpenStudy (anonymous):

take the log, get \[\frac{1}{x^2}\ln\left(\frac{\cos(x)}{\cos(2x)}\right)\] then maybe simplify using properties of the log, take the limit using l'hopital, and then exponentiate

OpenStudy (anonymous):

(cosx/cos 2x) ^ (1/x sqared). we started limits in calc but nothing with sin or cos. thats why im not sure how to even start this

OpenStudy (anonymous):

just started? do you know l'hopital's rule?

OpenStudy (anonymous):

not yet, he gave us this as a bonus. normally the limits we have to solve are for when x goes to + or - infinity and the problems are like [(2x^2+3x-4)/(3x^4-5x+6)]

OpenStudy (anonymous):

i cannot do this without l'hopital. i really have no idea what he expects. the limit inside the parentheses is 1 so you have the form \[1^{\infty}\] which is indetermined. you have to take the log, take the limit, and then exponentiate (that is, raise e to the power of whatever you get) to get your answer. i cannot think of an elementary way to do this. but the final answer will be \[e^{\frac{3}{2}}\] maybe i can ask someone for a simple method

OpenStudy (anonymous):

we have gone over: f(x)=e^(-1/x) and e^-infinity =0, and e^+infinity = +infinity. so would the limit be +infinity?

OpenStudy (jamesj):

Hm, if you don't use l'Hopital's rule, you're still going to need to use the derivative somewhere. How about this: do you know the power series expansion/Taylor series expansion/Maclaurien series expansion of cos x?

OpenStudy (anonymous):

in fact you need l'hopital twice. hard to imagine expanding in a taylor series before l'hopital

OpenStudy (anonymous):

i just thought maybe there was a gimmick i didn't see, but i am thinking there is not one

OpenStudy (jamesj):

(maybe. I learnt power series in high school before we discussed l'Hopital's rule.) I don't think there is one either.

OpenStudy (jamesj):

Good question though. But insanely hard if you don't have some of these sharp instruments in your tool box.

OpenStudy (anonymous):

So is my professor asking us to solve a problem he knows we wont be able to solve?....yet

OpenStudy (jamesj):

Perhaps he's doing it on purpose to help motivate the next thing he teaches you.

OpenStudy (anonymous):

thats what i was thinking, thanks for the help. and i just learned of this web site a few weeks ago and i think its fantastic. im not sure who to tell, but it deserves the praise!

OpenStudy (jamesj):

Glad you're finding it helpful.